OpenStudy (anonymous):

Evaluate the expression when a = 4 and b = 6. 8a – 4b + 5

OpenStudy (lieutenantgeneral):

you suppose to substitue a for for and b for 6

OpenStudy (lieutenantgeneral):

then you multiply 8*4-4*6+5

OpenStudy (anonymous):

8 x 4 - 4 x 6 + 5 = ?

OpenStudy (anonymous):

173?

OpenStudy (arabpride):

8a – 4b + 5 plug in the values: 8(4) - 4(6) + 5 = multiply: 8(4) - 4(6) + 5 = 32 - 24 + 5 8 + 5 = 13 your answer = 13

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

thanks (x

OpenStudy (anonymous):

8 x 4 - 4 x 6 + 5 = 13.

OpenStudy (arabpride):

No Problem *^.^*